OpenStudy (anonymous):

Solve equation using substitution x/3+y/5=7 x+y=27

OpenStudy (anonymous):

y=27-x replacing in the other equation: x/3 + (27-x)/5 =7 x/3 + 27/5 - x/5 =7 (5x-3x)/15 + 27/5 = 7 2x/15 + 27/5 = 7 2x/15 = 7 - 27/5 2x/15 = (35-27)/5 2x/15 = 8/5 2x = (15*8)/5 2x=24 x=12 then y = 27-x y=27-12 y=15